A string of length L, mass per unit length mu, and
Answer #1

Fundamental frequency of the vibrating string is f=\frac{1}{2L}\sqrt{\frac{T}{\mu}}

a) Length of the string is doubled.f_1=\frac{1}{2L_1}\sqrt{\frac{T}{\mu}}=\frac{1}{4L}\sqrt{\frac{T}{\mu}}=\frac{f}{2}

that is the frequency if halved.

b) Mass per unit length is doubled. f_1=\frac{1}{2L}\sqrt{\frac{T}{\mu_1}}=\frac{1}{2L}\sqrt{\frac{T}{2\mu}}=\frac{f}{\sqrt{2}}

c) Tension is doubled f_1=\frac{1}{2L}\sqrt{\frac{T_1}{\mu}}=\frac{1}{2L}\sqrt{\frac{2T}{\mu}}=\sqrt{2}f